OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 − 9x − (x + 12) less than or equal to 25? x greater than or equal to −3 x less than or equal to −3 −3 less than or equal to x A number line with a closed circle on negative 3 and shading to the right

jigglypuff314 (jigglypuff314):

Hello @asmith721 and sorry for the late reply :-/ in order to figure out which of them is NOT a way, we should first figure out what Are some ways to represent the given function? :) so we are given \(7-9x-(x+12)\le25\) we should first simplify this all down... (distributing the negative) \(7-9x - x - 12 \le 25 \) (adding like terms) (-9x - x = -10x) (7 - 12 = -5) \(-10x - 5 \le 25 \) (adding 5 to both sides) \(-10x \le 30\) (dividing both sides by -10) (but remember when you multiply or divide by a negative number, you should flip the sign) \(x \ge -3\) hopefully this would be enough to get you to your answer :) if you still need help, feel free to tag me with @jigglypuff314 ^_^